On January 1, 2016, you deposited $6,600 in a savings account. The account will earn 10 percent annual compound interest, which will be added to the fund balance at the end of each year.  

Required:
1.
What will be the balance in the savings account at the end of 4 years? (Future Value of $1, Present Value of $1, Future Value Annuity of $1, Present Value Annuity of $1.) (Use appropriate factor(s) from the tables provided. Round your final answers to 2 decimal places.)



2. What is the total interest for the 4 years? (Future Value of $1, Present Value of $1, Future Value Annuity of $1, Present Value Annuity of $1.) (Use appropriate factor(s) from the tables provided. Round your final answers to 2 decimal places.)



3. How much interest revenue did the fund earn in 2016 and in 2017? (Round your final answers to 2 decimal places.)
Answer #1

Solution 1:

Balance in saving account at the end of 4 years = P (1+i)^n = $6,600 * (1+0.10)^4 = $6,600*1.4641 = $9,663.06

Solution 2:

Total interest for the 4 years = $9,663.06 - $6,660 = $3,063.06

Solution 3:

Interest revenue for 2016 = $6,600* 10% = $660

Interest revenue for 2017 = ($6,600 + $660) * 10% = $726